21 E&K Company has a net book value of $250,000. The company has a 10 percent cost of capital. The firm expects to have profits of $45,000, $40,000, and $55,000 respectively for the next three years. The company pays no dividends and depreciation is five percent per year of book value.

  1. Using the excess earnings model, should the firm be valued at more or less than its book value?
  2. What, if anything, are the company’s abnormal earnings for the next three years?
  3. What is your best estimate of the firm’s value using the excess earnings model?
Answer #1

The firm shall be valued at more than the book value as the earnings are over and above the cost of capital.

Here, Cost of Capital = $25000

Depreciation = $12500

Earnings for all three years is above the cost of capital and depreciation combined.

The company's abnormal earnings for the next three years are:

Year 1 = 45000 - 37500 = $7500

Year 2 = 40000 - 37500 = $2500

Year 3 = 55000 - 37500 = $17500

Average excess earning = (7500 + 2500 + 17500)/3 = $9167

Cost of capital = 25000

Capitalised Value = (Cost of Capital + Average excess earning)/ cost of capital factor

= (25000+9167)/0.1

= $341667

Note: In the absence of information, the capitalisation method has been used to value the firm.